1.   A population consists of the numbers 2, 4, 9, 10, and 5. Let us list all possible samples size of 3 from this population. Find the mean and variance of the sampling distribution of the sample mean.


1
Expert's answer
2022-04-25T16:18:49-0400

The number of possible samples which can be selected without replacement is

"\begin{pmatrix}
 N \\
 n
\end{pmatrix}=\cfrac{N! } {n! \cdot(N-n)! }=\\
=\cfrac{5! } {3! \cdot2! }=\cfrac{4\cdot5}{2}=10."


All the possible samples of sizes n=3 wich can be drawn without replacement from the population:

"\{ (2,4,9), (2,4,10),(2,4,5),(2,9,10),(2,9,5),\\
(2,10,5),(4,9,10),(4,9,5),(4,10,5),(9,10,5)\}."


Population mean:

"\mu=\cfrac{2+4+9+10+5}{5}=6."


Population variance:

"\sigma^2=\sum(x_i-\mu)^2\cdot P(x_i),"

"X-\mu=\begin{Bmatrix}
 2-6,4-6,9-6,10-6,5-6
\end{Bmatrix}="

"=\begin{Bmatrix}
-4, - 2,3,4,-1
\end{Bmatrix},"

"\sigma^2=(-4)^2\cdot \cfrac{1}{5}+(-2)^2\cdot \cfrac{1}{5}+\\
+3^2\cdot \cfrac{1}{5}+4^2\cdot \cfrac{1}{5}+(-1)^2\cdot \cfrac{1}{5}=9.2."



Mean of the sampling distribution of sample means:

"\mu_{\bar x} =\mu=6."


Variance of the sampling distribution of sample means:

"\sigma^2_{\bar x}=\cfrac{\sigma^2}{n}=\cfrac{9.2}{3}=3.067."